Circular No. 6500

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bq IN NGC 996 C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discov-

ery of an apparent supernova on OCA films taken on Oct. 12.05 and 12.11 UT. SN 1996bq located is at α = 2^{h}38^{m}40^{s}.03, δ = +41^{o}39′17′′.8 (equinox 2000.0), which is 0′′.6 east and 27′′.4 north of the center of NGC 996. Nothing appears at this position on the original Palomar Sky Survey prints, on a second Sky Survey B film, or on 1995 OCA films. A recent deeper OCA film shows that the object has faded, as indicated by the following V magnitudes: Oct. 12.1, 18.7; Nov. 3.9, 19.4. A nearby star (V ∼ 18.5) has position end figures 40^{s}.07, 12′′.6.

COMETS C/1995 O1 (HALE-BOPP) AND 126P/IRAS D. Lynch, R. Russell, and T. Tessensohn, Aerospace Corporation; M.

Sitko, University of Cincinnati; and M. Hanner, Jet Propulsion Laboratory, report 8–13-μm spectroscopy of comet C/1995 O1 on Oct. 14.25 UT using the Infrared Telescope Facility (+ Aerospace BASS spectrograph) with a 3′′.4 beam and a 60′′ chop throw: “The comet’s central condensation showed a strong silicate emission feature with a line/continuum ratio of 1.5. The gross shape of the silicate emission closely resembles the now-familiar trapezoidal feature previously seen in comets 1P/Halley, C/1987 P1 (Bradfield), and C/1990 K1 (Levy). The feature also showed significant spectral structure with a peak at 11.2 ± 0.1 μm (0.4 μm wide, FWHM) and potentially a second peak at 10.1 ± 0.1 μm (0.3 μm wide). The 8–12.5-μm color temperature was 255 ± 15 K, or ∼ 85 K above the radiative equilibrium temperature (168 K) for a blackbody at heliocentric distance of 2.73 AU. Magnitudes at 8.0, 10.5, and 12 μm were 3.9 ± 0.1, 1.20 ± 0.06, and 1.16 ± 0.1, respectively. On Oct. 16.25, the spectrum was the same shape as before but ∼ 10 percent brighter. Photometry of comet 126P/IRAS was obtained on Oct. 15.3 with the same instrument; by averaging spectroscopic points between 10 and 13 μm, they obtained a positive detection of N′ = 2.7 ± 0.19 under poor seeing conditions.”

Visual m_{1} estimates of C/1995 O1: Oct. 8.39 UT, 5.0 (D. Seargent, The Entrance, N.S.W., naked eye); 13.74, 4.9 (M. Plsek, Lelekovice, Czech Rep., 10×50 binoculars); 29.74, 5.2 (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars); Nov. 1.76, 4.9 (R. Haver, Frasso Sabino, Italy, 10×50 binoculars); 3.71, 4.7 (V. Znojil, Brno, Czech Rep., naked eye).

1996 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green